OpenStudy (anonymous):

Solve the system of equations. 9x + y = 2 –4x – y = –17 A. (–3, 29) B. (1, –7) C. (0, 0) D. (4, 1)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Add the equations.

OpenStudy (anonymous):

I really don't know how to do this stuff but this is my last test so I will be done

OpenStudy (anonymous):

9x+y=2 -4x-y=-17 so 5x=-15 x=-3 pplug that back into one of the equations and solve for y

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you understand why?